# Industrial Internet of Things Intelligence Empowering Smart Manufacturing: A Literature Review

### 02 Dec 2023 v1

@@ VI-D Digital twin connects cyber-physical spaces
The digital twin attracts much attention in recent years, some great progresses are achieved. This paper reviews these contributions in terms of accurately modeling and interoperability.
In terms of accurately modeling in manufacturing, current works mainly focus on single-dimension modeling, including structure modeling, attribute modeling, behavior modeling, and process modeling. One finite element analysis approach orienting industrial rail surface wear is constructed [19]. The approach realizes fault monitoring of targets and the iteration of the finite element model by updating the key parameters. A digital twin model for atmospheric and vacuum petrochemical process is proposed [153], in order to achieve safety management and control of the petrochemical processes. For the assembly quality management in the manufacturing process, the surfaces of the assembly parts are modeled, and every stages of assembly process is tracked to reduce errors through online monitoring surface condition in the assembly process [20]. Driven by the increasingly complex and uncertain scheduling process of smart manufacturing system, the work [154] firstly introduces the digital twin modeling of the scheduling processes, and then illustrates some scheduling scheme generated by the algorithm by the twins. In this way, the effectiveness of production line scheduling is enhanced. However, researches of system-level (mixed/multiple-dimensions) modeling are lacking.
@@
Data-driven modeling, knowledge-based modeling, and hybrid modeling are the three most popular modeling methods for the digital twin. A neural network based modeling approach is proposed [155] for the uncertainty and complexity of the digital twin system. The digital twin technique is used into monitoring task of the tool Wear Monitoring of Micro-Milling [156]. The model is based on the existing electrical and mechanical formulas of the three core modules of the micro-milling machine tool. The work [157] presents a digital twin model that makes use of data collected by sensors, so that the performance (accuracy and reliability) can be improved by the knowledge and data analysis. A digital-twin-assisted fault diagnosis using deep transfer learning to analyze the operational conditions of machining tools is proposed[158]. This proposed system has developed an intelligent tool-holder that integrates a k-type thermocouple and cloud data acquisition system over the WiFi module, and the tool-holder can provide better accuracy to demonstrate the performance improvement of milling and drilling operations of cutting tools.
@@
Interoperability supports mutual execution between physical and digital spaces. It is an increasingly attractive topic. However, related researches are still at an early stage, and many problems remain about mutual intelligibility, interoperability, dynamic updating, data heterogeneity, submodel heterogeneity, etc. Currently, some works are achieved that focus more on solving heterogeneous data related problems. The work [159] studies the multi-scale evolution mechanism of product quality for the fusion and use of multi-source heterogeneous data. In the process, heterogeneous data is first classified and reorganized according to attributes and categories. Then according to the data level, it is divided into three categories: macro, medium, and micro. Finally, with the help of the dynamic knowledge base, the concepts of data mapping and rule mapping are used to form a scale dynamic quality map to describe the product quality. Jiang et al.[160] designed a data interaction mechanism that integrates message middleware, memory database and relational database for data interaction in virtual and physical space. The messaging middleware, eg Kafka and MQTT, is used to send production instructions and feedback execution results. The memory database, eg Redis, is used to store real-time operating data during production. The relational databases, eg MySQL, Oracle, SQL Server, and so on, are used for data management for simulation systems.
@@ VI-E Deep learning boosts intelligence
Deep learning methods use computational models that are composed of multiple processing layers to learn representations of data with multiple levels of abstraction [161]. It does not depend on prior data processing and automatically extracts features by using the backpropagation algorithm [162]. In recent years, deep convolutional nets have brought great breakthroughs in the domains of images, video, speech, and games, whereas recurrent neural networks do well in processing sequential data such as text and speech. Moreover, deep learning can speed up the development of IIoT intelligence.
Here deep learning methods are divided into three categories, supervised learning, unsupervised learning and deep reinforcement learning. Supervised learning infers a function from labeled training dataset where each training example is a pair that consists of an input object and a desired output value. Unsupervised learning must self-discover patterns from the training set in which any pre-assigned labels or scores are not provided. Reinforcement learning is an effective method that helps make a sequence of decisions to maximize the rewards by trial and error. This paper demonstrates the enabling effect of deep learning by researches in industrial machine vision as examples.
@@ VI-F Smart hardware builds intelligence foundation
Regarding the sensor technology, the complementary metal-oxide-semiconductor (CMOS) sensors have achieved commercial realization with satisfactory performance efficiency, replacing the charge-coupled devices (CCDs). Due to about 95% quantum efficiency, the latest CMOS sensors become nearly perfect detectors [185]. Event-based camera is a silicon retina which outputs asynchronous spikes with a number of benefits, including extremely low latency, high dynamic range, and very low power consumption [186][187]. For extremely low light imaging conditions, the quanta image sensor (QIS)—a single-photon detector—is worth noting for the unstructured industrial conditions [188][189]. Besides the optical sensors, computational imaging [190] is a wonderful breakthrough in industrial image process. It attempts to balance the processing capabilities of optics and electronics through concurrent design and joint optimization of all elements. It can extend the capability of conventional means, or reduce the measurement cost in comparison to conventional means. Specifically, the computational imaging contains a wide range of techniques, including the phase imaging, computed tomography (CT), magnetic resonance imaging (MRI), time-of-flight (ToF), light-field imaging, time-gated imaging, synthetic aperture radar, coded-aperture imaging, super-resolution, etc. These computational sensors have been adopted widely in industry [191, 192, 193, 194].
The computing capability keeps rapidly increasing due to Moore’s Law and other computing technologies (e.g. domain specific architectures) [195][196]. Therefore, the long data processing time required by computational imaging can be significantly shorter [197][198]. Besides, the platforms equipped with Intel CPU, Nvidia GPU, or Google TPU enable fast training of deep learning models [199][200].
VI-G Cloud/Edge computing inspires novel business and deployment
Cloud computing delivers computing services over the internet to users, and allows users to access and use resources such as computing power, storage, and software applications without the need for owning or managing physical infrastructure. In a cloud computing environment, services are provided on-demand, and users can scale their resources dynamically based on their requirements.
Edge computing is a distributed computing paradigm designed to process data in close proximity to the source of generation, minimizing latency and enhancing the efficiency of applications. Unlike traditional centralized cloud computing, edge computing involves local processing on devices or nearby servers. This proximity allows for real-time or near-real-time data processing, making it particularly suitable for applications such as Internet of Things (IoT) devices and augmented reality, where low latency is crucial.
There are many surveys discussing the technical models of cloud computing [201, 202, 203, 204] and edge computing [205, 206, 207]. This paper will not repeat it, but only focus on the contributions of cloud computing and edge computing in IIoT intelligence, that is, inspiring novel manufacturing business and deployment.
@@ VI-G1 Cloud manufacturing
Cloud manufacturing as a novel business mode can offer many advantages. Firstly, manufacturing resources, such as equipment, software tools, and computational power, are shared and accessed through the cloud [209, 210]. Advanced manufacturing resource virtualization and sharing strategies play a crucial role in optimizing production processes, improving efficiency, and making informed decisions [211, 212, 213]. This allows for more efficient utilization of resources and reduces the need for individual companies to invest heavily in specialized infrastructure. Secondly, manufacturers can access manufacturing services on a pay-as-you-go basis. This on-demand model allows companies to scale their production up or down based on market demand, without the need for significant capital investment. Thirdly, the cloud enables the collection and analysis of large volumes of data generated throughout the manufacturing process. Data analysis tools can be applied to gain insights into production efficiency, quality control, and overall performance. Finally, cloud manufacturing extends its impact beyond individual factories by integrating the entire supply chain. This integration improves communication, coordination, and responsiveness among suppliers, manufacturers, and distributors.
@@ VI-G2 Cloud-based deployment
Cloud-based deployment offers a wide range of benefits across various industries. Firstly, cloud computing eliminates the need for significant upfront investments in physical hardware and infrastructure. Enterprises can leverage a pay-as-you-go model, paying only for the computing resources they use. This results in cost savings, especially for small and medium-sized businesses that may not have the capital for extensive IT infrastructure. Secondly, cloud-based deployment allows manufacturers to scale their computing resources based on demand. This is particularly beneficial in smart manufacturing where data volumes can vary, and the need for computational power may fluctuate. Scalability ensures that resources are readily available to handle increased workloads during peak production times.. Thirdly, cloud-based deployment enable remote monitoring and management of manufacturing processes. This is particularly relevant in smart manufacturing, where sensors and IoT devices generate real-time data. Remote access allows stakeholders to monitor production, identify issues, and make informed decisions from anywhere with an internet connection. Finally, cloud computing provides an environment for manufacturing enterprises to experiment and adopt innovative technologies. This includes integrating artificial intelligence, machine learning, and other emerging technologies into smart manufacturing processes. Cloud-based platforms offer the computational power needed for training and deploying advanced models.
@@ VI-G3 Edge-based deployment
Edge-based deployment refers to the strategic placement of computing resources, data processing, and analysis capabilities closer to the source of data generation, typically at the network edge. Unlike cloud-based deployment where industrial data is sent to centralized servers for processing, edge-based deployment shifts some of the processing tasks to local devices or edge servers. This approach is particularly beneficial in industrial settings where real-time processing, low latency, and reliable connectivity are crucial for operational efficiency and decision-making [216, 217, 218].
Compared with cloud-based deployment, edge-based deployment can obtain advantages in low latency, bandwidth optimization, autonomous operation. Firstly, edge-based deployment processes data closer to the source reduces latency, enabling real-time or near-real-time responses. Secondly, edge-based deployment reduces the need to transmit large volumes of raw data to the cloud, optimizing bandwidth usage. This is especially important in scenarios where network bandwidth is limited or expensive. Thirdly, edge-Based deployment can operate autonomously even without continuous connectivity to the cloud. This ensures that critical functions can continue in the event of network disruptions, enhancing system reliability.
@@ VII-A Digital Control
It is challenging to realize the vision, but the digital twins connecting cyberspace and physical space can provide some opportunities after overcoming some difficulties. Firstly, due to the frequent interaction between digital twin models and industrial software to exchange real-time industrial data/commands, it is necessary to address interoperability issues between software and models, as well as standardization issues between communication protocols in industrial environments. Secondly, industrial datasets often suffer from issues such as noise, outliers, and missing values, impacting data quality, inconsistent data formats and units across different sources. Poor data quality leads to inaccurate analyses and unrealiable conclusions. Therefore, it is non-trivial to generate high-quality industrial data based on the practical data, in order to enable intelligent data analysis. Thirdly, at the beginning of digital twin design, due to our inability to collect/generate a perfect dataset that includes the behavior of modeled objects under all conditions, the performance of digital twins will decrease with the increase of invisible conditions. Therefore, it is necessary to design update mechanisms for digital twins to maintain the high-performance operation of the digital twin system.
@@ VII-C Cost-friendly operation and deployment
Cost-friendly deployment refers to the strategic and efficient implementation of systems or technologies with a focus on optimizing costs while achieving desired outcomes. The cost-friendly deployment of manufacturing processes is often realized by professionals who specialize in various aspects of industrial engineering, production management, and process optimization. These experts contribute their knowledge and skills to enhance efficiency, reduce costs, improve product quality, and streamline overall operations. We won’t interfere with the deployment optimization of manufacturing processes.
As the digital development of smart manufacturing, cost-friendly deployment of computation and storage resources is also required to ensure industrial data storage and analysis. Cloud-based deployment and edge-based deployment are able to provide cost-friendly deployment strategies for industrial big data and systems. However, based on the analysis in digital control and deterministic response, we can know that there are many time-critical tasks in IIoT. Both cloud-based deployment and edge-based deployment will introduce unexpected and unpredictable latency due to the usage of best-effort networks. Therefore, a novel local deployment strategy for computation and storage resources should be studied in the future, as well as the resource scheduling mechanisms, in order to ensure the deployment and scheduling strategies can satisfy latency requirement of all IIoT tasks within a constrained budget.
@@ VII-D IIoT intelligence proliferation
  • The potential scenarios for industrial manufacturing are diverse, and some situations may be unpredictable. Traditional manual design methods are difficult to fully cover all possible scenarios.
  • Intelligent proliferation can learn and understand patterns and regularities in industrial environments through accumulating industrial knowledge. When encountering emergency situations, intelligent proliferation system can automatically generate corresponding solutions to reduce the harm caused to the production process.
  • Intelligent emergence allows systems to automatically generate adaptive processing algorithms when facing new types of data tasks, significantly reducing the workload of independent development for each scenario. This adaptability can greatly enhance the system’s flexibility and efficiency.
@@
Inspired by the development of ChatGPT, which can understand and generate human-like text in conversations, retain context for coherent responses, perform tasks like answering questions and creative writing, and provide information based on its knowledge, a large industrial generative intelligent model has potential to achieve IIoT intelligence proliferation, and will be one of the future research directions.
VIII Conclusion
Due to the fiercely competitive business environment, increasingly personalized customization needs, and rising labor costs, the manufacturing industry is eager to undergo digital transformation. At the same time, many data-driven and knowledge-based intelligent technologies are emerging, lighting the way for intelligent manufacturing. IIoT intelligence refers to a series of techniques, methods, productions and platforms that can be adopted throughout the entire value chain to establish capabilities of digital connection and perception, intelligent analysis and cognition, immediate decisions, etc. It can provide innovative and efficient solutions for all aspects of the manufacturing value chain. Therefore, IIoT intelligence will drive the transformation and upgrading of smart manufacturing in terms of functionality, operation, deployment, and application.

### 22 Feb 2024 v2

@@ VI-D Digital twin connects cyber-physical spaces
The digital twin attracts much attention in recent years, and some great progresses have achieved. This paper reviews these contributions in terms of accurately modeling and interoperability.
@@
In terms of accurately modeling in manufacturing, current works mainly focus on single-dimension modeling, including structure modeling, attribute modeling, behavior modeling, and process modeling. Primary methods include data-driven modeling, knowledge-based modeling, and hybrid modeling. A neural network based modeling approach is proposed [146] for the uncertainty and complexity of the digital twin system. The digital twin technique is used into monitoring task of the tool Wear Monitoring of Micro-Milling [147]. The model is based on the existing electrical and mechanical formulas of the three core modules of the micro-milling machine tool. The work [148] presents a digital twin model that makes use of data collected by sensors, so that the performance (accuracy and reliability) can be improved by the knowledge and data analysis. A digital-twin-assisted fault diagnosis using deep transfer learning to analyze the operational conditions of machining tools is proposed[149]. This proposed system has developed an intelligent tool-holder that integrates a k-type thermocouple and cloud data acquisition system over the WiFi module, and the tool-holder can provide better accuracy to demonstrate the performance improvement of milling and drilling operations of cutting tools.
@@
Interoperability supports mutual execution between physical and digital spaces. It is an increasingly attractive topic. However, related researches are still at an early stage, and many problems remain about mutual intelligibility, interoperability, dynamic updating, data heterogeneity, submodel heterogeneity, etc. Currently, some works are achieved that focus more on solving heterogeneous data related problems. The work [150] studies the multi-scale evolution mechanism of product quality for the fusion and use of multi-source heterogeneous data. Jiang et al.[151] designed a data interaction mechanism that integrates message middleware, memory database and relational database for data interaction in virtual and physical space. The messaging middleware, such as Kafka and MQTT, is used to send production instructions and feedback execution results. The memory database, eg Redis, is used to store real-time operating data during production. The relational databases, eg MySQL, Oracle, SQL Server, and so on, are used for data management for simulation systems.
Although there are some challenges in the development of digital twin, its practical values cannot be ignored. According to the global lighthouse network 2023 [31], the digital twin has been used in many practical factories. For example, ACG Capsules in India has introduced digital twin to power production planning and scheduling, and improved 13% on-time delivery in full. Aramco in Saudi Arabia has developed a digital twin model for energy consumption reduction, and realized 159% profitability enhancement.
@@ VI-E Deep learning boosts intelligence
Deep learning methods use computational models that are composed of multiple processing layers to learn representations of data with multiple levels of abstraction [152]. It does not depend on prior data processing and automatically extracts features by using the backpropagation algorithm [153]. In recent years, deep learning has brought great breakthroughs [154, 155, 156, 157, 158] in the domains of scheduling and planning, computer vision, games, etc., and are speeding up the development of IIoT intelligence. Deep learning methods can be divided into three categories, supervised learning, unsupervised learning and deep reinforcement learning. This paper demonstrates the enabling effect of deep learning by researches in industrial machine vision as examples.
@@ VI-E3 Deep reinforcement learning
Currently, the practical intelligent transformation cannot leave deep learning. In Global Lighthouse Network 2023 [31], deep learning has been applied in every process and realize impacts everywhere. For example, Ingrasys deployed an AI-demand forecasting model, and the model has delivered 27% more accuracy; Hengtong Alpha Optic-Electric automatically optimizes perform and drawing parameters with a model trained on past strategies.
@@ VI-F Smart hardware builds intelligence foundation
The latest CMOS sensors, with approximately 95% quantum efficiency, have replaced CCDs due to their commercial realization and satisfactory performance efficiency in sensor technology [181]. Event-based cameras, known as silicon retinas, offer advantages such as low latency, high dynamic range, and low power consumption [182][183]. For low-light conditions, the quanta image sensor (QIS) serves as a valuable single-photon detector in unstructured industrial environments [184][185]. Besides the optical sensors, computational imaging [186] is another significant advancement in industrial image process. It optimizes the processing capabilities of optics and electronics through joint design and optimization. Techniques include phase imaging, CT, MRI, ToF imaging, light-field imaging, time-gated imaging, synthetic aperture radar, coded-aperture imaging, and super-resolution.These sensors find wide application across industries [187, 188, 189, 190].
The computing capability also keeps rapidly increasing due to Moore’s Law and other computing technologies (e.g. domain specific architectures) [191][192]. Therefore, the long data processing time required by computational imaging can be significantly shorter [193][194]. Besides, the platforms equipped with Intel CPU, Nvidia GPU, or Google TPU enable fast training of deep learning models [195][196].
VI-G Cloud/Edge computing inspires business and deployment
There are many surveys discussing the technical models of cloud computing [197, 198, 199, 200] and edge computing [201, 202, 203]. This paper will not repeat it, but focus on the contributions of cloud computing and edge computing in IIoT intelligence, including inspiring novel manufacturing business and deployment.
@@ VI-G1 Cloud manufacturing
Cloud manufacturing as a novel business mode can offer many advantages. Firstly, manufacturing resources, such as equipment, software tools, and computational power, are shared and accessed through the cloud [205, 206]. Advanced manufacturing resource virtualization and sharing strategies play a crucial role in optimizing production processes, improving efficiency, and making informed decisions [207, 208, 209]. This allows for more efficient utilization of resources and reduces the need for individual companies to invest heavily in specialized infrastructure. Secondly, manufacturers can access manufacturing services on a pay-as-you-go basis. This on-demand model allows companies to scale their production up or down based on market demand, without the need for significant capital investment. Thirdly, the cloud enables the collection and analysis of large volumes of data generated throughout the manufacturing process. Data analysis tools can be applied to gain insights into production efficiency, quality control, and overall performance.
@@ VI-G2 Cloud-based deployment
Cloud-based deployment offers a wide range of benefits across various industries. Firstly, cloud computing eliminates the need for significant upfront investments in physical hardware and infrastructure. Enterprises can leverage a pay-as-you-go model, paying only for the computing resources they use. This results in cost savings, especially for small and medium-sized businesses that may not have the capital for extensive IT infrastructure. Secondly, cloud-based deployment enable remote monitoring and management of manufacturing processes. This is particularly relevant in smart manufacturing, where sensors and IoT devices generate real-time data. Remote access allows stakeholders to monitor production, identify issues, and make informed decisions from anywhere with an internet connection. Thirdly, cloud computing provides an environment for manufacturing enterprises to experiment and adopt innovative technologies. This includes integrating artificial intelligence, machine learning, and other emerging technologies into smart manufacturing processes. Cloud-based platforms offer the computational power needed for training and deploying advanced models.
@@ VI-G3 Edge-based deployment
Unlike cloud-based deployment where industrial data is sent to centralized servers for processing, edge-based deployment shifts some of the processing tasks to local devices or edge servers. This approach is particularly beneficial in industrial settings where real-time processing, low latency, and reliable connectivity are crucial for operational efficiency and decision-making [212, 213, 214].
Edge-based deployment can obtain advantages in low latency, bandwidth optimization, autonomous operation. Firstly, edge-based deployment processes data closer to the source reduces latency, enabling real-time or near-real-time responses. Secondly, edge-based deployment reduces the need to transmit large volumes of raw data to the cloud, optimizing bandwidth usage. This is especially important in scenarios where network bandwidth is limited or expensive. Thirdly, edge-based deployment can operate autonomously even without continuous connectivity to the cloud. This ensures that critical functions can continue in the event of network disruptions, enhancing system reliability.
In practice, multiple cloud companies are cooperating with manufacturing to realize business novelty and deployment optimization. For example, by utilizing Microsoft Azure high-performance computing and other Microsoft technologies, STMicroelectronics is advancing its research and development, transforming its supply chain, scaling manufacturing, and giving its employees the right tools to secure its position as a world leader in semiconductor solutions [215]. Phillips chose to modernize its legacy on-premises Microsoft Dynamics solution by upgrading to the latest cloud capability of Microsoft Dynamics 365 and Dynamics 365 Field Service. The upgrade provides the company with real-time insights into the performance of the business, which has helped Phillips make process improvements and helped increase operational efficiency and accuracy of field resources [216].
@@ VII-A Digital control
Although achieving digital control poses challenges, the digital twins connecting cyberspace and physical space offer opportunities once three key difficulties are addressed:
  • Overcoming issues between software and models, as well as standardization challenges in communication protocols within industrial environments, is crucial.
  • Ensuring high-quality industrial data is essential for accurate analyses and reliable conclusions. Generating such data from practical datasets is vital for enabling intelligent data analysis.
  • Building digital twins requires datasets covering the behavior of modeled objects under various conditions. However, it’s impractical to collect perfect datasets for all scenarios. Thus, implementing update mechanisms is necessary to maintain high-performance operation of the digital twin system.
@@ VII-B Deterministic response
Building an innovative computing and networks convergence paradigm with the capability of deterministic responses will be helpful. Some researchers have started to explore such a paradigm [217, 218, 219, 220], but some problems are waiting to be deeply studied.
  • How to schedule existing IIoT computation and network resources to ensure deterministic responses for scalable computing tasks?
  • How to design sharing mechanisms that support deterministic scheduling of both communication traffics and computing tasks on the same IIoT networks?
@@ VII-C Cost-friendly operation and deployment
Cost-friendly deployment refers to the strategic and efficient implementation of systems or technologies with a focus on optimizing costs while achieving desired outcomes. The cost-friendly deployment of manufacturing processes is often realized by professionals who specialize in various aspects of industrial engineering, production management, and process optimization. We won’t interfere with the deployment optimization of manufacturing processes, but consider the cost-friendly deployment of computation and storage resources to ensure industrial data storage and analysis. Cloud-based deployment and edge-based deployment are able to provide cost-friendly deployment strategies for industrial big data and systems. However, there are many time-critical tasks in IIoT. Both cloud-based deployment and edge-based deployment can lead to failures in meeting the time requirements of tasks due to unexpected and unpredictable latency introduced by the usage of best-effort networks. Therefore, following two challenges should be studied in the future to build cost-friendly deployment.
  • Given the personalized deployment costs associated with computing servers and network devices, how to design a cost-effective deployment solution for IIoT enterprises that lack a foundation for intelligent transformation?
  • For IIoT enterprises with established intelligent transformation foundations, how can the current deployment structure of computing servers and network devices be optimized to cater to the personalized transformation needs?
@@ VII-D IIoT intelligence proliferation
Inspired by the development of ChatGPT, which can understand and generate human-like text in conversations, retain context for coherent responses, perform tasks like answering questions and creative writing, and provide information based on its knowledge, a large industrial generative intelligent model has potential to achieve IIoT intelligence proliferation, but it comes with several challenges:
  • Building and maintaining large models can be costly, especially in terms of computational resources and data storage.
  • Large models demand extensive and high-quality training data, which may be scarce or challenging to obtain in industrial settings.
  • Deploying large models in real-world industrial environments can be complex due to resource constraints and real-time processing requirements.
VIII Ethical Implications and Environmental Impacts
In this section, we will explore the ethical implications and environmental impacts of adopting Industrial Internet of Things (IIoT) intelligence in manufacturing.
VIII-A Ethical implications
The ethical implications primarily include three aspects: job displacement and workforce transition, privacy and data security, social and economic equality.
VIII-A1 Job displacement and workforce transition
The increased automation facilitated by IIoT intelligence may lead to job displacement for certain manual tasks, raising ethical concerns about the well-being and transition of affected workers. To avoid the unexpected situations, manufacturing enterprises that will realize transformation with IIoT intelligence are expected to carry out three steps [221]. Before the introduction of new technologies, companies should communicate the benefits and explain the why, make the bigger picture clear to workers of the technology, help workers to explore and become confident with the technologies, and include workers in the exchange of ideas. During the introduction of new technology, companies should make timelines clear early on, and provide a diversified skill training programmes to workers. After the introduction of new technology, companies should follow up the progress of technology applications, and encourage workers together to continue to explore new use cases.
VIII-A2 Privacy and data security
IIoT intelligence involves vast operations of data collection and analysis, raising privacy concerns for employees and stakeholders. To address this concern, companies should be transparent about their data collection, processing, and usage practices in IIoT deployments. This transparency helps build trust with employees, stakeholders, and the public. Furthermore, detailed privacy policies and data usage guidelines should be communicated clearly to all relevant parties, outlining the types of data collected, the purposes for which it is used, and the rights of individuals regarding their data. Additionally, encryption techniques should be employed to secure data both in transit and at rest, ensuring that it remains confidential and tamper-proof.
VIII-A3 Social and economic equality
Larger corporations with greater financial resources and technical capabilities may have an advantage in adopting and leveraging IIoT solutions, leaving smaller manufacturers or those in less developed regions at a disadvantage. This unequal access can exacerbate existing disparities in productivity, competitiveness, and innovation within the manufacturing sector. To avoid the unequal access of IIoT intelligence, IIoT solution providers are expected to offer flexible and scalable offerings tailored to the needs and budget constraints of manufacturing businesses of all sizes. The development of open-source IIoT platforms and standards should be encouraged to facilitate interoperability and lowers barriers to entry for manufacturers of all scales. Initiatives such as establishing innovation clusters, technology parks, or shared manufacturing facilities can be established to encourage collaboration, knowledge sharing, and access to advanced technologies among local manufacturers.
VIII-B Environmental Impacts
Adopting IIoT intelligence in manufacturing can be both positive and negative.
VIII-B1 Positive impacts
IIoT intelligence enables better monitoring and optimization of resource usage, such as energy, water, and raw materials, leading to reduced waste and improved resource efficiency. Optimizing manufacturing processes through IIoT intelligence can lead to reduced emissions of greenhouse gases and other pollutants. IIoT-enabled predictive maintenance and quality control measures reduce the likelihood of defects and production errors, resulting in less waste and scrap materials. Additionally, IIoT intelligence provides greater visibility into supply chains, enabling more efficient logistics, inventory management, and transportation planning.
VIII-B2 Negative impacts
The proliferation of IIoT devices and sensors may contribute to a significant increase in electronic waste at the end of their lifecycle. While IIoT technologies can improve energy efficiency in manufacturing, the deployment of additional sensors, data centers, and network infrastructure may lead to increased overall energy consumption, especially, the deep application of the industrial large model will consume vase amount of electronic energy. The production of IIoT devices, including sensors, processors, and communication modules, requires the extraction of raw materials and energy-intensive manufacturing processes.
VIII-B3 Mitigation strategies
To reduce the side effects of IIoT intelligence on environment, some mitigation strategies are suggested. Firstly, IIoT devices for durability, repairability, and recyclability are encouraged to be designed to promote a circular economy approach, minimizing waste and resource consumption. Secondly, engaging suppliers committed to sustainable development, purchase environmentally friendly materials, reduce emissions related to transportation, and establish a more sustainable IIoT supply chain. Thirdly, conducting comprehensive lifecycle assessments and environmental impact analyses helps identify areas for improvement and inform decision-making throughout the IIoT deployment process.
IX Conclusion